# SPIdepth: Strengthened Pose Information for Self-supervised Monocular Depth Estimation

### v2

@@ Abstract
Self-supervised monocular depth estimation has garnered considerable attention for its applications in autonomous driving and robotics. While recent methods have made strides in leveraging techniques like the Self Query Layer (SQL) to infer depth from motion, they often overlook the potential of strengthening pose information. In this paper, we introduce SPIdepth, a novel approach that prioritizes enhancing the pose network for improved depth estimation. Building upon the foundation laid by SQL, SPIdepth emphasizes the importance of pose information in capturing fine-grained scene structures. By enhancing the pose network’s capabilities, SPIdepth achieves remarkable advancements in scene understanding and depth estimation. Experimental results on benchmark datasets such as KITTI, Cityscapes, and Make3D showcase SPIdepth’s state-of-the-art performance, surpassing previous methods by significant margins. Specifically, SPIdepth tops the self-supervised KITTI benchmark. Additionally, SPIdepth achieves the lowest AbsRel (0.029), SqRel (0.069), and RMSE (1.394) on KITTI, establishing new state-of-the-art results. On Cityscapes, SPIdepth shows improvements over SQLdepth of 21.7% in AbsRel, 36.8% in SqRel, and 16.5% in RMSE, even without using motion masks. On Make3D, SPIdepth in zero-shot outperforms all other models. Remarkably, SPIdepth achieves these results using only a single image for inference, surpassing even methods that utilize video sequences for inference, thus demonstrating its efficacy and efficiency in real-world applications. Our approach represents a significant leap forward in self-supervised monocular depth estimation, underscoring the importance of strengthening pose information for advancing scene understanding in real-world applications.
@@ 4 Results
Our assessment of SMIDepth encompasses three widely-used datasets: KITT, Cityscapes and Make3D, employing established evaluation metrics.
@@ 4.1.2 Cityscapes Dataset
Cityscapes [6] poses a unique challenge with its plethora of dynamic objects. To gauge SMIDepth’s adaptability, we fine-tune on Cityscapes using pre-trained models from KITTI. Notably, we abstain from leveraging motion masks, a feature common among other methods, even in the presence of dynamic objects. Our performance improvements hinge solely on SMIDepth’s design and generalization capacity. This approach allows us to scrutinize SMIDepth’s robustness in dynamic environments. We adhere to data preprocessing practices from [61], ensuring consistency by preprocessing image sequences into triples.
@@ 4.1.3 Make3D Dataset
Make3D [38] is a monocular depth estimation dataset containing 400 high-resolution RGB and low-resolution depth map pairs for training, and 134 test samples. To evaluate SMIDepth’s generalization ability on unseen data, zero-shot evaluation on the Make3D test set has been performed using the SMIDepth model pre-trained on KITTI.
@@ 4.2 KITTI Results
We present the performance comparison of SMIDepth with several state-of-the-art self-supervised depth estimation models on the KITTI dataset, as summarized in Table 1. SMIDepth achieves superior performance compared to all other models across various evaluation metrics. Notably, it achieves the lowest values of AbsRel (0.071), SqRel (0.531), RMSE (3.662), and RMSElog (0.153), indicating its exceptional accuracy in predicting depth values.
Moving on to Table 2, we compare the performance of SMIDepth with several supervised depth estimation models on the KITTI eigen benchmark. Despite being self-supervised and metric fine-tuned, SMIDepth outperforms supervised methods across all these metrics, indicating its superior accuracy in predicting metric depth values.
Furthermore, SMIDepth surpasses LightedDepth, a model that operates on video sequences (more than one frame) and outperforms a good pre-trained models like EVP based on stable diffusion [36]. Despite LightedDepth’s advantage of using multiple frames, SMIDepth shows improvements of 0.012 (29.3%) in AbsRel, 0.038 (34.3%) in SqRel, 0.354 (20.3%) in RMSE, and 0.011 (18.6%) in RMSElog, highlighting SMIDepth’s robustness and effectiveness even in challenging scenarios.
Additionally, SMIDepth demonstrates significant performance improvements over SQLdepth, a model that serves as the foundation for its development. In the self-supervised setting, SMIDepth shows improvements of 5.3% in AbsRel, 1.5% in SqRel, 1.6% in RMSE, and 1.9% in RMSElog. In the supervised setting, SMIDepth shows improvements of 32.6% in AbsRel, 35.6% in SqRel, 17.9% in RMSE, and 25% in RMSElog. These substantial improvements underscore the impact of strengthening the pose net and its information in SMIDepth.
Overall, these results underscore the effectiveness of SMIDepth in self-supervised monocular depth estimation, positioning it as a leading model in the field. Qualitative results further illustrate the superior performance of SMIDepth, as shown in Figure 2.
@@ 4.3 Cityscapes Results
To evaluate the generalization of SMIDepth, we conducted fine-tuning experiments in a self-supervised manner without using a motion mask on the Cityscapes dataset. Starting from a KITTI pre-trained model, we fine-tuned it on Cityscapes. The results, summarized in Table 3, demonstrate that SMIDepth outperforms all other methods, including those that use motion masks.
Despite not using a motion mask—a technique commonly employed to handle the high proportion of moving objects in the Cityscapes dataset—SMIDepth achieves remarkable improvements over other models. Compared to SQLdepth, SMIDepth shows significant advancements: improvements of 0.023 (21.7%) in AbsRel, 0.432 (36.8%) in SqRel, and 1.032 (16.5%) in RMSE.
Moreover, compared to the previous state-of-the-art model RM-Depth, which also uses motion masks, SMIDepth achieves improvements of 0.007 (7.8%) in AbsRel, 0.084 (10.2%) in SqRel, and 0.298 (5.4%) in RMSE.
These results underscore SMIDepth’s exceptional generalization and accuracy, achieved without the use of motion masks. This makes SMIDepth a highly robust and efficient option for depth estimation tasks. Its performance demonstrates its capability for quick deployment in new datasets, effectively addressing the challenges posed by moving objects.
@@ 4.4 Make3D results
To assess the generalization capacity of SMIDepth, a zero-shot evaluation was performed on the Make3D dataset [38] using pretrained weights from KITTI. Adhering to the evaluation settings of [15, 45], SMIDepth achieved superior results compared to other methods, including SQLdepth. Table 4 highlights these findings, showcasing the remarkable zero-shot generalization ability of the SMIDepth model.
As summarized in Table 4, SMIDepth achieves the lowest values in all evaluation metrics, with AbsRel (0.299), SqRel (1.931), RMSE (6.672), and $log_{10}$ (0.144). These results highlight the remarkable zero-shot generalization ability of the SMIDepth model, significantly outperforming the previous best model, SQLdepth. The improvements of SMIDepth over SQLdepth are 0.007 (2.3%) in AbsRel, 0.471 (19.6%) in SqRel, 0.184 (2.7%) in RMSE, and 0.007 (4.6%) in $log_{10}$, underscoring its superior performance in challenging zero-shot scenarios.
@@ 6 Conclusion
These results highlight SPIdepth’s potential for real-world applications, offering precise depth estimation and superior performance across diverse challenges, ready for rapid adaptation and deployment.

### v3

@@ Abstract
Self-supervised monocular depth estimation has garnered considerable attention for its applications in autonomous driving and robotics. While recent methods have made strides in leveraging techniques like the Self Query Layer (SQL) to infer depth from motion, they often overlook the potential of strengthening pose information. In this paper, we introduce SPIdepth, a novel approach that prioritizes enhancing the pose network for improved depth estimation. Building upon the foundation laid by SQL, SPIdepth emphasizes the importance of pose information in capturing fine-grained scene structures. By enhancing the pose network’s capabilities, SPIdepth achieves remarkable advancements in scene understanding and depth estimation. Experimental results on benchmark datasets such as KITTI, Cityscapes, and Make3D showcase SPIdepth’s state-of-the-art performance, surpassing previous methods by significant margins. Specifically, SPIdepth tops the self-supervised KITTI benchmark. Additionally, SPIdepth achieves the lowest AbsRel (0.029), SqRel (0.069), and RMSE (1.394) on KITTI, establishing new state-of-the-art results. On Cityscapes, SPIdepth shows improvements over SQLdepth of 21.7% in AbsRel, 36.8% in SqRel, and 16.5% in RMSE, even without using motion masks. On Make3D, SPIdepth in zero-shot outperforms all other models. Remarkably, SPIdepth achieves these results using only a single image for inference, surpassing even methods that utilize video sequences for inference, thus demonstrating its efficacy and efficiency in real-world applications. Our approach represents a significant leap forward in self-supervised monocular depth estimation, underscoring the importance of strengthening pose information for advancing scene understanding in real-world applications. The code and pre-trained models are publicly available at https://github.com/Lavreniuk/SPIdepth.
@@ 4 Results
Our assessment of SPIDepth encompasses three widely-used datasets: KITTI, Cityscapes and Make3D, employing established evaluation metrics.
@@ 4.1.2 Cityscapes Dataset
Cityscapes [6] poses a unique challenge with its plethora of dynamic objects. To gauge SPIDepth’s adaptability, we fine-tune on Cityscapes using pre-trained models from KITTI. Notably, we abstain from leveraging motion masks, a feature common among other methods, even in the presence of dynamic objects. Our performance improvements hinge solely on SPIDepth’s design and generalization capacity. This approach allows us to scrutinize SPIDepth’s robustness in dynamic environments. We adhere to data preprocessing practices from [61], ensuring consistency by preprocessing image sequences into triples.
@@ 4.1.3 Make3D Dataset
Make3D [38] is a monocular depth estimation dataset containing 400 high-resolution RGB and low-resolution depth map pairs for training, and 134 test samples. To evaluate SPIDepth’s generalization ability on unseen data, zero-shot evaluation on the Make3D test set has been performed using the SPIDepth model pre-trained on KITTI.
@@ 4.2 KITTI Results
We present the performance comparison of SPIDepth with several state-of-the-art self-supervised depth estimation models on the KITTI dataset, as summarized in Table 1. SPIDepth achieves superior performance compared to all other models across various evaluation metrics. Notably, it achieves the lowest values of AbsRel (0.071), SqRel (0.531), RMSE (3.662), and RMSElog (0.153), indicating its exceptional accuracy in predicting depth values.
Moving on to Table 2, we compare the performance of SPIDepth with several supervised depth estimation models on the KITTI eigen benchmark. Despite being self-supervised and metric fine-tuned, SPIDepth outperforms supervised methods across all these metrics, indicating its superior accuracy in predicting metric depth values.
Furthermore, SPIDepth surpasses LightedDepth, a model that operates on video sequences (more than one frame) and outperforms a good pre-trained models like EVP based on stable diffusion [36]. Despite LightedDepth’s advantage of using multiple frames, SPIDepth shows improvements of 0.012 (29.3%) in AbsRel, 0.038 (34.3%) in SqRel, 0.354 (20.3%) in RMSE, and 0.011 (18.6%) in RMSElog, highlighting SPIDepth’s robustness and effectiveness even in challenging scenarios.
Additionally, SPIDepth demonstrates significant performance improvements over SQLdepth, a model that serves as the foundation for its development. In the self-supervised setting, SPIDepth shows improvements of 5.3% in AbsRel, 1.5% in SqRel, 1.6% in RMSE, and 1.9% in RMSElog. In the supervised setting, SPIDepth shows improvements of 32.6% in AbsRel, 35.6% in SqRel, 17.9% in RMSE, and 25% in RMSElog. These substantial improvements underscore the impact of strengthening the pose net and its information in SPIDepth.
Overall, these results underscore the effectiveness of SPIDepth in self-supervised monocular depth estimation, positioning it as a leading model in the field. Qualitative results further illustrate the superior performance of SPIDepth, as shown in Figure 2.
@@ 4.3 Cityscapes Results
To evaluate the generalization of SPIDepth, we conducted fine-tuning experiments in a self-supervised manner without using a motion mask on the Cityscapes dataset. Starting from a KITTI pre-trained model, we fine-tuned it on Cityscapes. The results, summarized in Table 3, demonstrate that SPIDepth outperforms all other methods, including those that use motion masks.
Despite not using a motion mask—a technique commonly employed to handle the high proportion of moving objects in the Cityscapes dataset—SPIDepth achieves remarkable improvements over other models. Compared to SQLdepth, SPIDepth shows significant advancements: improvements of 0.023 (21.7%) in AbsRel, 0.432 (36.8%) in SqRel, and 1.032 (16.5%) in RMSE.
Moreover, compared to the previous state-of-the-art model RM-Depth, which also uses motion masks, SPIDepth achieves improvements of 0.007 (7.8%) in AbsRel, 0.084 (10.2%) in SqRel, and 0.298 (5.4%) in RMSE.
These results underscore SPIDepth’s exceptional generalization and accuracy, achieved without the use of motion masks. This makes SPIDepth a highly robust and efficient option for depth estimation tasks. Its performance demonstrates its capability for quick deployment in new datasets, effectively addressing the challenges posed by moving objects.
@@ 4.4 Make3D results
To assess the generalization capacity of SPIDepth, a zero-shot evaluation was performed on the Make3D dataset [38] using pretrained weights from KITTI. Adhering to the evaluation settings of [15, 45], SPIDepth achieved superior results compared to other methods, including SQLdepth. Table 4 highlights these findings, showcasing the remarkable zero-shot generalization ability of the SPIDepth model.
As summarized in Table 4, SPIDepth achieves the lowest values in all evaluation metrics, with AbsRel (0.299), SqRel (1.931), RMSE (6.672), and $log_{10}$ (0.144). These results highlight the remarkable zero-shot generalization ability of the SPIDepth model, significantly outperforming the previous best model, SQLdepth. The improvements of SPIDepth over SQLdepth are 0.007 (2.3%) in AbsRel, 0.471 (19.6%) in SqRel, 0.184 (2.7%) in RMSE, and 0.007 (4.6%) in $log_{10}$, underscoring its superior performance in challenging zero-shot scenarios.
@@ 6 Conclusion
These results highlight SPIdepth’s potential for real-world applications, offering precise depth estimation and superior performance across diverse challenges. Its lightweight design and adaptability make it an ideal candidate for integration into various systems, enabling rapid deployment and scalable solutions in environments where accurate depth perception is crucial.